# Shuttling Compiler for Trapped-Ion Quantum Computers Based on Large Language Models

### v1

@@ Abstract
Trapped-ion quantum computers based on segmented traps rely on shuttling operations to establish connectivity between multiple sub-registers within a quantum processing unit. Several architectures of increasing complexity have already been realized, including linear arrays, racetrack loops, and junction-based layouts. As hardware capabilities advance, the need arises for flexible software layers within the control stack to manage qubit routing—the process of dynamically reconfiguring qubit positions so that all qubits involved in a gate operation are co-located within the same segment. Existing approaches typically employ architecture-specific heuristics, which become impractical as system complexity grows. To address this challenge, we propose a layout-independent compilation strategy based on large language models (LLMs). Specifically, we fine-tune pretrained LLMs to generate the required shuttling operations. We evaluate this approach on both linear and branched one-dimensional architectures, demonstrating that it provides a foundation for developing LLM-based shuttling compilers for trapped-ion quantum computers.
@@ 1 Introduction
Shuttling-based trapped-ion quantum computers [KielpinskiMW02, KausahlLSHPSBMSP20, PinoDFGMABFHMRN21, MuraliDBM20, MosesB+23] are currently limited by both qubit count and the depth of reliably executable quantum circuits. Recent advances are gradually overcoming these barriers, enabling demonstrations of quantum advantage [AruteAB+19] and opening pathways toward practical applications in real-world computational tasks [PeruzzoMSYZLAO14, AnschuetzOAC19]. At the same time, this progress introduces additional hardware complexity, necessitating more sophisticated compilation strategies for efficient qubit routing and quantum circuit execution.
@@
In trapped-ion systems, atomic ions are confined in segmented microchip traps [SchulzPSS06]. Each pair of electrodes generates a localized potential well that can hold multiple ions, effectively partitioning the trap into smaller, more manageable segments. The quantum information is encoded in the internal states of individual ions, each of which represents a qubit; accordingly, we use the term qubit to refer to an ion throughout the paper. By adjusting electrode voltages, trapping potentials can be dynamically reconfigured, enabling qubits to be shuttled between segments for successive gate operations. Segments are typically categorized as either storage segments, where qubits are held, or gate segments, where quantum operations are performed using laser or microwave radiation. In accordance with the current maturity level of trapped-ion hardware, the number of available gate segments is still very limited and often restricted to a single one [KausahlLSHPSBMSP20, SchmaleTBPKDOWB22].
@@
Most current architectures arrange trap segments in a linear array [KausahlLSHPSBMSP20, PinoDFGMABFHMRN21, BowlerGLTHJHLW12, LeeJPJKC21] or in a racetrack loop [MosesB+23], where each segment connects to two neighbors. Such designs limit scalability, as the number of qubits scales linearly with the physical size of the trap. Introducing junctions [HensingerOSHYADMR06, BlakestadOVABLW09, WrightAFVDHPLDKSH13, ShuVBNVSB14] increases connectivity, allowing segments to connect to multiple neighbors. Quantinuum has demonstrated this idea in their Helios architecture, where a ring-shaped storage region is connected to two linear legs via an $\mathsf{X}$-shaped junction [Ransford+25]. However, Helios contains only a single junction. Increasing the number of junctions would enable more complex topologies in the future, such as two-dimensional grids [LekitschWFMDWH17].
@@
Executing a quantum circuit on such hardware requires a sequence of physical operations. Because quantum gates can only be applied to qubits within the same gate segment, the trap must be dynamically reconfigured using four fundamental shuttling operations:
  • Translation: Moves the qubits from one trap segment to a neighboring segment.
@@
  • Separation: Splits the qubits stored in one segment into two segments.
  • Merge: Combines qubits from two segments into a single segment.
@@
  • Swap: Physically rotates the content of a segment by 180 degrees, reversing the order of the qubits within the segment.
@@
While shuttling operations enable effective all-to-all connectivity between qubits and strongly suppress cross-talk errors, they also introduce substantial time overhead that lengthens operation timescales [HilderPOSOLRMSP22]. As a result, error sources such as dephasing and noise-induced heating become more pronounced, while the shuttling process itself can further degrade gate fidelities through ion heating. The gate fidelity quantifies how closely the implemented quantum gate matches its ideal operation, with lower fidelities indicating higher error probabilities during gate execution. Consequently, the number of shuttling operations must be minimized, highlighting the need for efficient compilation routines.
@@
Compilation refers to the translation of a logical quantum circuit, which comprise abstract qubits and idealized gates, into a sequence of hardware-level instructions executable on a specific trapped-ion device. This process includes decomposing logical gates into the device’s native gate set, mapping logical to physical qubits, and determining a shuttling schedule that defines an executable sequence of shuttling operations. The resulting schedule must adhere to hardware constraints while minimizing resource overheads such as operation time, qubit transport distance, and cumulative error.
Existing shuttling compilers compute valid qubit routing schedules by generating shuttling schedules tailored to specific architectures, such as linear arrays [PinoDFGMABFHMRN21, Wagner22, DurandauWMBSPB23], racetrack loops [MosesB+23], drive-through architectures [ChangJCHL25], plus-sign layouts [SchmaleTBPKDOWB22], a ring with linear legs [Ransford+25], branched one-dimensional traps [KreppelMWHPSB24], and two-dimensional designs [WebberHWH20, DaiBR24, SchoenbergerHBW25, SchoenbergerHSW25, SchoenbergerW25]. However, as architectures grow more complex, designing separate heuristics for each layout becomes increasingly impractical.
@@
To overcome the limitations of architecture-specific compilation, large language models (LLMs) offer a promising, layout-independent alternative. LLMs can perform new tasks from textual instructions or a few demonstrations [Brown+20], making them well suited for learning shuttling behavior directly from examples.
@@
In this work, we investigate whether current LLMs can generate valid and efficient shuttling schedules and identify the challenges that arise when applying them to trapped-ion shuttling. We fine-tune LLaMA 3.2 [MetaAI24], Qwen 3 [QwenTeam25], DeepSeek LLM [Deepseek-AI24], and Gemma 3 [GemmaTeam25] on datasets derived from classical shuttling compilers for both linear [Wagner22, DurandauWMBSPB23] and branched one-dimensional architectures with junctions [KreppelMWHPSB24]. We then compare their performance across multiple benchmark circuits and additionally evaluate their ability to generalize to previously unseen trap layouts.
Our evaluation shows that the fine-tuned models were able to generate valid shuttling schedules and, in some cases, even surpassed classical shuttling compilers. However, performance degraded as the number of qubits and gates increased: all models frequently reached trap configurations that require numerous retries or prevent further schedule generation. Results on the unseen trap layouts further indicate that the models’ ability to generalize beyond the training architectures remains limited. Overall, the results show that our approach provides a solid foundation for developing LLM-based shuttling compilers.
@@
Our compiler differs from machine-learning-based routing techniques developed for superconducting architectures [PalerSFA23, PozziHSM22, FanGL22, PascoalFA24, TangDKFKS24, SinhaAS22, RussoPPAC25, SangHH25, ZeynaliB25, CuomoCKTAPC23, EscofetOBPVFAAA25, SundaramGR25]. While those methods primarily optimize SWAP placement and gate parallelism on fixed coupling graphs, our approach accounts for the hardware-specific characteristics of trapped-ion devices. In particular, it considers shuttling operations, which are absent in superconducting systems, and the limited number of gate segments, which restrict the parallel execution of quantum gates. Our work is the first to employ large language models for compiling quantum circuits into shuttling schedules for trapped-ion quantum computers and to compare different LLMs for this task.
@@ 2 Related Work
Within these compilation stacks, the individual compilation stages can be realized using a variety of algorithmic techniques. Compilation typically begins with a transformation to the hardware’s native gate set. Different methods pursue different optimization goals, such as reducing circuit depth, minimizing the number of specific gates, or adapting to particular architectures. Some compilers apply heuristics to obtain approximate solutions [YounisI22, JinHCHZZ23, KreppelMOWHPSB23, AroraXWLLDA25], while others employ SAT solvers to produce optimal results for smaller instances [PehamBKWB23]. More recently, machine learning has been applied to this task, ranging from recommender models [WeidenYKKI23] to reinforcement learning agents trained for specific optimization goals [MoroPRP21, FöselNML+22, Preti24, RietschDUPPMS24, Ruiz+25, RiuNVGE25, WeidenKKI25].
After this transformation, logical qubits must be mapped to physical qubits, and routing must be computed to enable the execution of gates under architectural constraints. This step is architecture-dependent, and techniques have been developed for a variety of platforms. For example, methods exist for neutral atom architectures [SchmidPW24, NottinghamPSWBCB24, StadeSVW24] and spin-qubit systems [ParaskevopoulosSAF23, ParaskevopoulosAF24], but most work focuses on superconducting and trapped-ion platforms.
In superconducting devices, architectures are modeled as coupling graphs, where vertices represent qubits and edges denote possible two-qubit interactions. Since two-qubit gates are restricted to adjacent qubits, additional SWAP gates are required to establish long-range connectivity. At the same time, gates on independent qubits can be executed simultaneously. Existing approaches differ in their optimization goals, such as minimizing SWAPs or maximizing gate parallelism. Proposed techniques span a wide range of methods, including heuristics and mathematical formulations [MaslovFM08, SaeediWD11, ShafaeiSP14, WilleLD14, LinSJ15, PedramS16, SiraichiSCP18, HattoriY18, LiDX19, MuraliBJCM19, ChildsSU19, CowtanDDKSS19, MatsuoY19, ZulehnerBW19, ZulehnerPW19, ItokoRIM20, NiuSST20, AlamAG20, TanC20, ZhouLF20, ZhangHQJYZ21, LiWSJDX22, ParkKKSK22, NanniciniBGJ22, HuaJCVKBLJZ23, PehamBW23, KhandavilliPNLND23, LiuYWHKI23, FuZWXSL23, ShaikP23, YangKSHD24, ChengYKWCH24, KattemölleH25], as well as machine-learning-based approaches. The latter include k-nearest neighbors, random forests, and multilayer perceptrons for qubit mapping [PalerSFA23], and reinforcement learning (RL) methods for routing [PozziHSM22, FanGL22, PascoalFA24, TangDKFKS24]. Specifically, [TangDKFKS24] employs a transformer-based neural network as an RL agent and reports a reduction in the number of SWAP gates by 10 to 20 % compared to the SABRE heuristic [LiDX19]. Hybrid approaches combining reinforcement learning with graph neural networks (GNNs) have also been proposed [SinhaAS22].
Beyond single-device architectures, several works address distributed superconducting architectures, where qubits are partitioned across multiple devices and minimizing inter-device qubit transfers becomes critical [RussoPPAC25, SangHH25, ZeynaliB25, CuomoCKTAPC23, EscofetOBPVFAAA25, SundaramGR25]. In this setting, GNN- and transformer-based models are used to capture global circuit structure [RussoPPAC25, SangHH25], and qubit reuse strategies are explored to further reduce communication overhead [SangHH25]. These approaches demonstrate reductions of up to 100 % in inter-device qubit transfers, effectively eliminating them when qubit reuse is enabled, and reductions of up to 99 % compared to QUBO [BandicPNORAVVAAF23] even without qubit reuse. The improvements are attributed to RL-based methods learning long-range dependencies across the entire circuit, enabling globally optimized qubit placements. Compared to Qiskit [WilleMN19], however, the reduction in inter-device transfers is in the range of 15 to 40 %, and some circuits still incur higher communication costs. Finally, [ZeynaliB25] proposes a noise-adaptive mapping strategy for distributed architectures that combines reinforcement learning for qubit placement with long short-term memory (LSTM) networks for noise migration and a multi-head self-attention mechanism to identify long-range qubit dependencies. Experimental results show that this approach increases the mean circuit fidelity by 49 %, raising it from 0.62 to 0.92, while reducing inter-chip operations by an average of 80 % compared to QUBO.
@@
Trapped-ion quantum computers exhibit distinct hardware characteristics. In these systems, multiple qubits are confined together within individual trap segments. While quantum gates can only be applied at specific segments, the qubit distribution across segments can be dynamically reconfigured through shuttling operations. Shuttling compilers compute sequences of such operations, enabling circuit execution on a given trap. Starting from a quantum circuit, they map logical qubits to physical ones and then derive shuttling schedules. While the minimal number of timesteps can be computed using Boolean satisfiability methods [SchoenbergerHBW24], most existing compilers employ heuristics to reduce the number of shuttling operations. Prior work includes design studies analyzing how architectural choices affect shuttling overhead [MuraliDBM20, AshSakiTG22], as well as compilers targeting specific trap configurations as mentioned in Sec. 1. However, all these existing approaches are tightly coupled to specific architectures.
@@
Beyond single-trap systems, compilation has also been explored for distributed trapped-ion architectures, where multiple traps are connected via photonic interconnects that mediate quantum communication between nodes [WuZWW25]. Such systems introduce additional scheduling and routing challenges, further highlighting the need for flexible compilation strategies that generalize across architectures.
In contrast to prior work, we present the first layout-independent shuttling compiler for trapped-ion quantum computers based on large language models (LLMs) and provide a comparison of different models for this task. By fine-tuning on examples of shuttling problems, LLMs are able to generalize beyond a single trap design and generate valid and efficient shuttling schedules for a variety of architectures, provided appropriate training data is available.
@@ 3 Compiler
This section explains how we fine-tune LLMs to generate shuttling schedules for quantum circuit execution. The process comprises three phases, illustrated in Fig. 1 and detailed in the following subsections.
@@ 3.1 Dataset Generation
To fine-tune an LLM, we require examples that illustrate how shuttling is performed. These examples are provided as shuttling schedules, as shown in Fig. 2. Since the physical movements depend only on which qubits participate in each gate, rather than on the specific gate function, the gate functionality is omitted in this representation. Such schedules can be generated in different ways, for example by constructing random sequences, by execution guided through a reward function, or by employing another (not necessarily optimized) compiler. The number of required schedules depends on the trap architecture, its size, and the number of qubits. Larger and more complex traps, as well as higher qubit counts, demand more shuttling operations per gate. This increases both the difficulty for the LLM to learn generating valid shuttling schedules and the amount of data required for fine-tuning.
@@
Each data entry follows the Alpaca format [TaoriGZDLGLH23], consisting of an instruction and an output. Since the LLM can only process textual input, all required information is encoded as text.
@@ 3.1.1 Instruction
Fig. 4: Quantum circuit with four qubits and 14 gates. Purple gates form the first layer and are executable when their qubits are located in a gate segment, whereas yellow gates belong to deeper layers and are not yet executable. For each qubit, gates are numbered in ascending order to indicate their execution sequence.
In addition, the instruction specifies the quantum gates of the circuit, which must be executed on the given architecture. Each gate targets specific qubits and is assigned a unique natural number for identification. Due to architectural constraints, a gate is executable only if all involved qubits reside in the same gate segment, and this segment must not contain any additional qubits. As illustrated in Fig. 4, the gates are divided into two groups: gates in the first layer of the circuit which are executable when the corresponding qubits are in a gate segment and not yet executable gates in deeper layers. For each qubit, the gate numbers are assigned in ascending order. To ensure the gate execution in the correct order, gates targeting the same qubit must be executed in numerical order, while gates acting on different qubits may be executed in any order. Finally, to link the circuit qubits with the qubits in the trap, all qubits in the gate descriptions are replaced by tuples $[v,p]$ of their corresponding position in the trap.
@@
In segmented trapped-ion architectures, junctions are formed by the intersections of electrodes, enabling qubits to be routed between different linear sections. Unlike storage segments, junctions do not provide a stable trapping potential suitable for holding qubits. Instead, the electrode potentials are shaped to guide qubits smoothly through the junction, ensuring efficient transport. Consequently, qubits must pass through junctions without stopping, and reversing direction within a junction is infeasible because the electric fields are engineered for unidirectional traversal rather than oscillatory motion [WrightAFVDHPLDKSH13, MokhberiSW17]. When computing shuttling operations, the LLM must respect this physical constraint. In particular, Separate, Merge, Swap, and gate execution cannot be performed on junction vertices, which must remain empty throughout execution. Furthermore, for Separate and Merge, the adjacent vertices involved must not be junctions. The operations are formally defined as follows:
@@
  • Execute Gate: Perform one of the first-layer gates in a gate segment, provided all previously defined execution conditions are satisfied.
@@
    First-layer gates: All gates in the first layer of the circuit that are executable when their corresponding qubits reside in a gate segment, along with the tuples of the qubits they act on.
@@
The instruction concludes by defining the task for the LLM: generate a shuttling sequence that enables the execution of one of the first-layer gates, contributing to the overall goal. After each operation, the LLM should also provide the updated qubit positions, the list of first-layer gates with updated position tuples of the qubits, and the new set of allowed shuttling operations. Requiring these additional outputs at each step supports the LLM to reason carefully about the current state and the implications of each operation.

### v2

@@ Abstract
Trapped-ion quantum computers based on segmented traps rely on shuttling operations to establish long-range connectivity between sub-registers. Qubit routing dynamically reconfigures qubit positions so that all qubits involved in a gate operation are co-located within the same segment, a task whose complexity increases with system size. To address this challenge, we propose a layout-independent compilation strategy based on large language models (LLMs). Specifically, we fine-tune pretrained LLMs to generate the required shuttling operations. We evaluate this approach on linear and branched one-dimensional architectures using quantum circuits of up to 16 qubits. Our results show that the fine-tuned LLMs generate valid shuttling schedules and, in some cases, outperform previous shuttling compilers by requiring approximately 15 % less shuttle overhead. However, results degrade as the algorithms increase in width and depth. In future, we plan to improve LLM-based shuttle compilation by enhancing our training pipeline using Direct Preference Optimization (DPO) and Gradient Regularized Policy Optimization (GRPO).
@@ 1 Introduction
Shuttling-based trapped-ion quantum computers [28, 27, 47, 41, 40] are currently limited by the number of qubits and the quality of gate operations. We are addressing the complexity that arises when a large number of qubits is routed efficiently for the execution of quantum algorithms.
@@
For a scalable shuttle-based approach of a quantum computer, the quantum information is encoded in the internal states of individual ions, each of which represents a qubit; accordingly, we use the term qubit to refer to an ion throughout the paper. The qubits are confined in segmented microchip traps [62], where each segment generates a localized potential well. By adjusting electrode voltages, trapping potentials can be dynamically reconfigured, enabling qubits to be shuttled between segments. Segments are either storage segments, where qubits are held, or gate segments. When two qubits are in the same gate segment, laser- or microwave-driven gate operations are executed. In accordance with the current maturity level of trapped-ion hardware, the number of available gate segments is still very limited [27, 57]. In this work we restrict ourselves to a single gate segment, while the approach can be extended to a number of gate segments, which have been realized recently [40, 52].
@@
Current architectures arrange trap segments in a linear array [27, 47, 5, 33] or in a racetrack loop [40], where each segment connects to two neighbors. Such designs limit scalability, as the number of qubits scales linearly with the physical size of the trap. Introducing junctions [24, 4, 75, 64] increases connectivity, allowing segments to connect to multiple neighbors. Quantinuum has demonstrated this idea in their Helios architecture, where a ring-shaped storage region is connected to two linear legs via an $\mathsf{X}$-shaped junction [52]. However, Helios contains only a single junction. Increasing the number of junctions would enable more complex topologies in the future, such as a two-dimensional quantum CCD [34].
@@
Executing a quantum circuit requires a sequence of physical operations. The trap must be dynamically reconfigured using four fundamental shuttling operations:
  • Translation: Moves the qubits to a neighboring segment.
@@
  • Separation: Splits two qubits from one segment into two different segments.
  • Merge: Joins one qubit from the left and one from the right into one segment.
@@
  • Swap: Physically rotates within one segment, thus reversing the order of the qubits.
@@
Shuttling operations introduce substantial time overhead [25]. As a result, error sources such as dephasing and noise-induced heating become more pronounced, and the shuttling process itself can further degrade gate fidelities through ion motional excitation and heating. Consequently, the number of shuttling operations must be minimized, highlighting the need for efficient compilation routines.
@@
Compilation refers to the translation of a quantum circuit, which comprises abstract qubits and idealized gates, into a sequence of instructions executable on a specific trapped-ion device. This process starts out by decomposing gates into the device’s native gate set, then mapping qubits to trapped ions, and finally determining a shuttling schedule that defines an executable sequence of shuttling operations. The resulting schedule must adhere to the architecture of the processor while minimizing resource overheads such as operation time, qubit transport distance, etc.
Existing shuttling compilers compute valid qubit routing schedules by generating shuttling schedules tailored to specific architectures [47, 71, 19, 40, 7, 57, 52, 30, 72, 14, 59, 58, 61]. Due to the architectures growing complexity, designing separate heuristics for each layout becomes increasingly impractical.
@@
We investigate large language models (LLMs) as a layout-independent alternative. LLMs can perform new tasks from textual instructions or a few demonstrations [6], making them well suited for learning shuttling behavior directly from examples.
@@
In this work, we examine whether current LLMs can generate valid and efficient shuttling schedules and identify the challenges that arise when applying them to trapped-ion shuttling. We fine-tune LLaMA 3.2 [37], Qwen 3 [49], DeepSeek LLM [16], and Gemma 3 [22] on datasets derived from classical shuttling compilers for both linear [71, 19] and branched one-dimensional architectures with junctions [30]. We then compare the performance across multiple benchmark circuits. Moreover, we evaluate the ability to generalize to novel trap layouts.
@@
Our compiler differs from machine-learning-based routing techniques developed for superconducting architectures [45, 48, 21, 46, 68, 65, 55, 56, 77, 13, 20, 67]. While those methods primarily optimize SWAP placement and gate parallelism on fixed coupling graphs, our approach accounts for the hardware-specific characteristics of trapped-ion devices. Our work is the first to employ LLMs for compiling quantum circuits into shuttling schedules for trapped-ion quantum computers and to compare different LLMs for this task.
@@ 2 Related Work
Due to the hardware characteristics of shuttling-based trapped-ion systems, shuttling compilers are a crucial component of the quantum compilation stack. Starting from a quantum circuit, they map qubits to ions and then derive shuttling schedules. While the minimal number of timesteps can be computed using Boolean satisfiability methods [60], most existing compilers employ heuristics to reduce the number of shuttling operations. Prior work includes design studies analyzing how architectural choices affect shuttling overhead [41, 1], as well as compilers targeting specific trap configurations [47, 71, 19, 40, 7, 57, 52, 30, 72, 14, 59, 58, 61]. However, all these existing approaches are tightly coupled to specific architectures.
@@
Beyond single-trap systems, compilation has also been explored for distributed trapped-ion architectures, where multiple traps are connected via photonic interconnects that mediate quantum communication between nodes [76]. Such systems introduce additional scheduling and routing challenges, further highlighting the need for flexible compilation strategies that generalize across architectures.
@@ 3 Compiler
This section explains how we fine-tune LLMs to generate shuttling schedules for quantum circuit execution. The process comprises three phases: dataset generation, fine-tuning, and inference.
@@ 3.1 Dataset Generation
First, we require examples that illustrate how shuttling is performed. These examples are provided as shuttling schedules, as shown in Fig. 1. Since the physical movements depend only on which qubits participate in each gate, rather than on the specific gate function, the gate functionality is omitted in this representation. Schedules can be generated in different ways, for example by constructing random sequences, by execution guided through a reward function, or by employing another (not necessarily optimized) compiler. The number of required schedules depends on the trap architecture, its size, and the number of qubits. Larger and more complex traps, as well as higher qubit counts, demand more shuttling operations per gate. This increases both the difficulty for the LLM to learn generating valid shuttling schedules and the amount of data required for fine-tuning.
@@
Each data entry follows the Alpaca format [69], consisting of an instruction and an output. Since the LLM can only process textual input, all required information is encoded as text.
@@ 3.1.1 Instruction
In addition, the instruction specifies the quantum gates of the circuit, which must be executed on the given architecture. Each gate targets specific qubits and is assigned a unique natural number for identification. Due to architectural constraints, a gate is executable only if all involved qubits reside in the same gate segment, and this segment must not contain any additional qubits. The quantum circuit can be partitioned into layers such that the first layer contains only gates whose involved qubits have no predecessors and are executable once the corresponding qubits are in the gate segment, while gates in deeper layers have at least one qubit with a predecessor and are therefore not yet executable. This layer-based division is illustrated in Fig. 3. For each qubit, the gate numbers are assigned in ascending order. To ensure the gate execution in the correct order, gates targeting the same qubit must be executed in numerical order, while gates acting on different qubits may be executed in any order. Finally, to link the circuit qubits with the qubits in the trap, all qubits in the gate descriptions are replaced by tuples $[v,p]$ of their corresponding position in the trap.
Fig. 3: Quantum circuit with four qubits and 14 gates. Purple gates form the first layer and are executable when their qubits are located in the gate segment, whereas yellow gates belong to deeper layers and are not yet executable. For each qubit, gates are numbered in ascending order to indicate their execution sequence.
@@
In segmented trapped-ion architectures, junctions are formed by the intersections of electrodes, enabling qubits to be routed between different linear sections. Qubits must pass through junctions without stopping, and reversing direction within a junction is infeasible because the electric fields are engineered for unidirectional traversal rather than oscillatory motion [75, 39]. In particular, Separate, Merge, Swap, and gate execution cannot be performed on junction vertices, which must remain empty throughout execution. Furthermore, for Separate and Merge, the adjacent vertices involved must not be junctions. The operations are formally defined as follows:
@@
  • Execute Gate: Perform the quantum gate in the gate segment, once all previously defined execution conditions are satisfied.
@@
    First-layer gates: All gates in the first layer of the circuit that are executable when their corresponding qubits reside in the gate segment, along with the tuples of the qubits they act on.
@@
The instruction concludes by defining the task for the LLM: generate a shuttling sequence that enables the execution of one of the first-layer gates, contributing to the overall goal. After each operation, the LLM should also provide the updated qubit positions, the list of first-layer gates with the updated position tuples of the qubits, and the new set of allowed shuttling operations. Requiring these additional outputs at each step supports the LLM to reason carefully about the current state and the implications of each operation.